### v1

# thanks: Equal contributionsthanks: Equal contributions

Abstract
Viscous films flowing down an incline can form self-sustained running waves, known as Kapitsa roll waves. Here we describe an analogous electron-hydrodynamic instability that produces similar running waves in Dirac materials such as graphene mono- and multilayers. It arises when carrier kinetics near charge neutrality make current dissipation strongly density-dependent. As the flow velocity $u$ exceeds a critical value, the system transitions to a state with coupled spatial and temporal oscillations. Experimentally, the instability should manifest as (i) a nonanalytic behavior characteristic of a second-order transition—an abrupt increase in time-averaged current—and (ii) narrow-band emission at the characteristic “washboard” frequency $f=u/\lambda$, where $\lambda$ is the modulation wavelength. This behavior parallels the AC and DC transport of sliding charge-density waves, but here it originates from a distinct, intrinsic mechanism unrelated to disorder. Estimates indicate that the emission frequency $f$, tunable by current, spans a broad range, highlighting Dirac bands as a promising platform for high-frequency electron-fluid dynamics.
Pattern formation is a hallmark of driven many-body systems, revealing how collective behavior can emerge even without external structuring. A paradigmatic example is the Turing instability [1, 2, 3]. Despite its universality, this mechanism has been explored almost exclusively in classical settings. Here we predict a quantum analogue: a modulational (Turing-type) instability in quantum materials driven out of equilibrium by a DC current. The key idea is that an applied current selectively softens the system at particular wave numbers, triggering an instability toward spontaneous spatial modulation. We find that current-driven Turing effects naturally arise in quantum dissipative systems, such as Dirac and Weyl metals in which electron-hole scattering enhances resistivity near charge neutrality while remaining weak away from it [4, 5, 6, 7, 8, 9, 10, 11].
Beyond fundamental interest, determining when a current-carrying state becomes unstable to self-oscillation may also clarify a longstanding practical question: Can a DC current generate coherent high-frequency radiation? Proposed mechanisms have ranged from Bloch oscillations [12, 13, 14, 15] to plasmonic instabilities [16, 17], and while experimental signatures exist [18, 19, 20, 21, 22], achieving robust and tunable self-oscillation has remained difficult. Our results identify a new route to current-driven pattern formation and self-oscillation— arising from the interplay of dissipation, band geometry, and hydrodynamic effects in Dirac materials — and highlight fresh opportunities for exploring quantum systems with complex band geometry or strong correlations.
Figure 1: (a) Current dissipation in graphene bilayer (BLG) band vs. electron density. Dissipation is high in the Dirac plasma near charge neutrality point (CNP) but drops rapidly upon doping away from CNP. This behavior can drive the Turing–Kapitsa instability. Insets illustrate carrier population in BLG band for densities at CNP and away from it, respectively. (b) Instability growth rate ${\rm Im}\,\omega_{\pm}(k)$ for the modes $\omega_{+}(k)$ and $\omega_{-}(k)$, Eq. (18), representing density and velocity perturbations. The density mode satisfies $\omega_{+}(0)=0$ by particle-number conservation, becoming unstable above a critical flow velocity (Eq.(2)); the growth rate ${\rm Im}\,\omega_{+}(k)>0$ peaks at the wave numbers $\pm k_{*}$ marked by arrows. The velocity mode remains damped, ${\rm Im}\,\omega_{-}(k)<0$. Small initial perturbations grow into a periodic modulation with wavelength $\lambda\sim 2\pi/k_{*}$ propagating downstream, as shown in Fig. 2.
The instability mechanism proposed here is reminiscent of the Kapitsa instability in viscous films flowing down an incline [32, 33, 34, 35, 36, 37]. The emergence of periodic modulations in steady laminar flows of open-surface films is a direct consequence of the balance between gravity and viscous drag: $\rho g\sin\theta=\eta\nabla^{2}{\boldsymbol{v}}({\boldsymbol{r}})$, where $\theta$ is the incline angle, $g$ gravity, and $\eta$ viscosity. A steady state defines a scaling between flow velocity and film thickness, $u\propto w^{2}g/\eta$. Since $u$ grows rapidly with $w$, regions with slightly larger film thickness $w(x)$ will flow faster, thereby amplifying perturbations and giving rise to a self-sustained running wave,
$w(x)=\bar{w}+\delta w\cos k(x-ut).$ (1)
In fluid films, the wavelength $\lambda=2\pi/k$ and modulation amplitude are set by the competition between viscous and capillary forces [33]. The emerging wavenumber $k$ coincides with $k_{*}$—the wavevector at which the instability growth rate is maximal—mirroring the essence of Turing’s pattern-formation mechanism [34].
As we will see, an electronic analogue of this instability can cause a steady (DC) electron flow to spontaneously develop self-sustained charge and current oscillations. The correspondence between an atomically thin 2D conductor with carrier density $n(x,y)$ and a classical viscous film with spatially varying thickness w(x,y)w(x,y)—measured between the substrate and its free surface—is established by identifying $n(x,y)$ with $w(x,y)$. Beyond this “hyperspace mapping” $w(x,y)\to(x,y,n)$, realizing the Turing–Kapitsa instability requires a dissipation mechanism whose rate, being strong, varies rapidly with nn—the electronic analogue of a drag coefficient in viscous films that decreases sharply with increasing film thickness.
Figure 2: Wavelength selection at instability. Shown is the long-time modulation (orange) evolved from different initial perturbations (blue): $\delta n(x)=\sum_{m}a_{m}\cos k_{m}x+b_{m}\sin k_{m}x$, $\delta p(x)=\sum_{m}a^{\prime}_{m}\cos k_{m}x+b^{\prime}_{m}\sin k_{m}x$, with $k_{m}=\frac{2\pi}{L}\times 3,5,8,13,21,55$. Irrespective of initial conditions, the system converges to the same spatial pattern, with wavelength set by the wavenumber $k$ of maximal growth (marked by arrows in Fig. 1), consistent with Turing’s maximum-growth rule.
There are several mechanisms reported in the literature that cause dissipation to decrease as carrier concentration increases. Here we focus on current dissipation in Dirac bands due to electron–hole (e–h) scattering at charge neutrality (CNP) [4, 5, 6, 7, 8, 9, 10, 11]. This dissipation channel is strong at CNP, where the Dirac plasma is e–h balanced, but it weakens rapidly once the carrier density is tuned away from neutrality. The resulting density-dependent dissipation plays a role similar to the width-dependent viscous drag in thin-film hydrodynamics: when the drift velocity exceeds a threshold, the flow becomes unstable and develops downstream-propagating waves. Our analysis predicts a critical drift velocity (see below):
$u_{c}=\frac{u_{0}}{R},\qquad R=\frac{n}{\gamma}\left|d\gamma/dn\right|,\qquad u_{0}=\sqrt{Un/m},$ (2)
where the dimensionless quantity $R$ characterizes how sensitively the dissipation rate $\gamma$ responds to changes in carrier density; $u_{0}$ is 2D plasmon velocity at density $n$. We emphasize that, because the dissipative regime of interest lies close to CNP, the relevant carrier densities are much smaller—i.e., much closer to neutrality—than those typically employed in studies of ballistic 2D plasmon propagation [38, 39, 40]. As illustrated in Fig. 1(a), the strong $T$-dependence of $\gamma$ leads to a pronounced $n$-dependence of $R$ at finite carrier densities near CNP. As discussed in [41], this causes the critical velocity $u_{c}$ to vary strongly with both $T$ and $n$.
The self-sustained modulation that appears for $u>u_{c}$ follows Turing’s wavelength-selection rule. Near threshold, the modulation wavelength is $\lambda=2\pi/k_{*}$, where $k_{*}$ is the wavenumber of the most unstable harmonic—i.e., the mode at which the instability growth rate is maximal (see Figs. 1(b) and 2). The transition is continuous and second-order (see Fig. 3). For realistic parameters, the resulting periodicity $\lambda$ can vary widely, from submicron to hundreds of microns [41].
Observable signatures of this collective dynamics resemble those studied for other running wave phenomena, such as sliding charge-density and spin-density waves (CDW and SDW) [42, 43, 44]. One is a narrow-band emission at the “washboard” frequency $f=u/\lambda$, where $u$ is the flow velocity tunable by the applied DC current. Another is a characteristic nonlinearity in the current–field dependence at the instability threshold, shown in Fig.3. The increase in the slope of the current vs. $u$ dependence reflects the contribution of running waves to nonlinear conductivity, resembling the Lee-Fukuyama-Rice conductivity of sliding CDW[42, 43, 44].
We stress that both the AC emission by running waves at a frequency $f=u/\lambda$ and the DC current nonlinearity are intrinsic effects, present already in a homogeneous system. This is in contrast to CDW sliding, where narrow-band emission and current nonlinearity are governed by the effects due to pinning-depinning by disorder.
Figure 3: Signatures and observables for current-driven electronic Turing–Kapitsa instability. Plotted are the running waves modulation amplitude and time-averaged current near the instability onset, obtained numerically for a graphene bilayer model (see Eqs. (10), (11) and accompanying text). The red line shows the theoretical square-root behavior $\Delta n\propto(u_{d}-u_{d}^{*})^{1/2}$ expected on the general symmetry grounds [52] The sharp slope change in the time-averaged (DC) current–field dependence marks the instability threshold and the added current from the running waves. Insets (a) and (b) illustrate a typical modulation and the current–field relation.
To explore the emergence of running waves under Turing-Kapitsa instability driven by a DC current, we consider a simple but realistic transport model. A natural framework to analyze collective dynamics of electrons is provided by hydrodynamic equations for particle number and momentum density $n(x,t)$ and ${\boldsymbol{p}}(x,t)$. Microscopically, these quantities are
$n(x,t)=\sum_{{\boldsymbol{p}}^{\prime}}f({\boldsymbol{p}}^{\prime},x,t),\quad{\boldsymbol{p}}(x,t)=\sum_{{\boldsymbol{p}}^{\prime}}{\boldsymbol{p}}^{\prime}f({\boldsymbol{p}}^{\prime},x,t),$ (3)
where $f({\boldsymbol{p}},x,t)$ is particle phase-space distribution summed over spins and valleys that evolves in time under the kinetic equation [45, 46]
$(\partial_{t}+{\boldsymbol{v}}\cdot\partial_{{\boldsymbol{x}}}+{\boldsymbol{F}}\cdot\partial_{{\boldsymbol{p}}})f({\boldsymbol{p}},\xi)=I[f].$ (4)
Here, ${\boldsymbol{v}}=\partial_{{\boldsymbol{p}}}\epsilon(p)$ is the velocity, ${\boldsymbol{F}}=e{\boldsymbol{E}}$ the force, and $I[f]$ the collision integral describing electron-electron and electron-hole scattering, and interactions with disorder and phonons; $\xi=(x,t)$ denotes space and time variables.
We are interested in transport in a Dirac band of a graphene bilayer, $\varepsilon_{\pm}(\mathbf{k})=\pm\tfrac{1}{2}\!\left(\sqrt{\gamma_{1}^{2}+4(\hbar vk)^{2}}-\gamma_{1}\right)$, where $v\approx 10^{6}\,$m/s and $\gamma_{1}=0.39\,$eV. For simplicity, we consider a one-species model representing one component of the bilayer bands—either electron or hole. The results below are largely insensitive to the exact form of $\epsilon(p)$ or to microscopic details of how momentum dissipation varies with carrier density. We model this by a dissipation rate that peaks in the region of width $w\sim k_{B}T$ and decreases with detuning from the CNP[10, 11]:
$\gamma(n)=\gamma_{0}e^{-|n|/w},\quad w=\frac{2m}{\hbar^{2}}k_{B}T,\quad m\approx 0.04m_{\rm e},$ (5)
where $w$ describes the CNP width due to intrinsic e-h scattering in pristine graphene in a parabolic band approximation. In this model, the electron temperature is replaced by an effective time-averaged value, neglecting its dynamics. Although heating is important, it occurs on time scales much longer than those of oscillatory electron motion. Moreover, recent experiments[47, 48, 49] show that high currents can flow through moiré graphene at drift velocities approaching the band velocity without significant overheating. Efficient cooling via the hBN substrate enables such high-current operation without damaging the graphene or broadening the Dirac band at CNP.
When applied to the particle number and momentum densities, $n$ and $p$, Eq.(4) gives conservation laws:
$\left(\gamma(n)+\partial_{t}\right)$ ${\boldsymbol{p}}(\xi)+\partial_{{\boldsymbol{x}}}\Pi_{0}(\xi)=en\left({\boldsymbol{E}}-\partial_{{\boldsymbol{x}}}\Phi(\xi)\right)$
$\partial_{t}$ $n(\xi)+\partial_{{\boldsymbol{x}}}[n(\xi){\boldsymbol{v}}(\xi)]=0.$ (6)
Here the first equation is the continuum version of the second Newton’s law $dp/dt=F$, the second equation is the continuity equation for $n$. In a parabolic band, momentum density is related to particle current density ${\boldsymbol{j}}=\sum_{p^{\prime}}{\boldsymbol{v}}(p^{\prime})f({\boldsymbol{p}}^{\prime})\equiv n{\boldsymbol{v}}$ as $m{\boldsymbol{j}}={\boldsymbol{p}}$.
The quantity $\Pi_{0}$ is the Fermi pressure of a flowing Fermi sea (FS). For Doppler-shifted Fermi distribution $f({\boldsymbol{p}})=\frac{1}{e^{\beta(\epsilon(p)-{\boldsymbol{u}}\cdot{\boldsymbol{p}}-\mu)}+1}$ describing electron gas of density $n$ flowing with the velocity $u$. For parabolic band, $\Pi_{0}$ is a sum of a static contribution and a kinetic energy term
$\Pi_{0}=\sum_{p^{\prime}}p_{x}^{\prime}v_{x}(p^{\prime})f({\boldsymbol{p}}^{\prime})=\frac{n^{2}}{2\nu}+\frac{p^{2}}{nm},\quad p=mnu,$ (7)
where $\nu=n/\epsilon_{F}=gm/2\pi\hbar^{2}$ is the density of states ($g$ is spin and valley degeneracy). The simple form of the result in Eq.(7), which is a sum of the pressure of a Fermi sea at rest $\Pi_{0}(n)=\sum_{p^{\prime}}\epsilon(p^{\prime})f_{u=0}(p^{\prime})=n^{2}/2\nu$ and the Bernoulli pressure $p^{2}/nm$ due to the flow, is a property of a parabolic band arising from its Galilean symmetry.
The transport equations in Eqs.(6) capture the essential physics underlying the current-driven Turing instability. A more realistic description, however, must explicitly include the interaction and viscous effects incorporated in the hydrodynamic stress density $\Pi$.
Below, for conciseness, we use a constant interaction model deferring the discussion of long-range interaction till later. In this case, the last term in (6) takes the form
$\Phi(\xi)=U_{0}[n(\xi)-\bar{n}],$ (8)
where constant $\bar{n}$ represents an offset charge on a gate.
It is convenient to combine the terms $\partial_{{\boldsymbol{x}}}\Pi_{0}$ and $\partial_{{\boldsymbol{x}}}\Phi$ in Eq. (6) into a net Fermi pressure term $\partial_{{\boldsymbol{x}}}\Pi$ that accounts for both electron–electron interactions and the flow, [50]
$\Pi=\frac{1}{2}U\left(n(\xi)\right)^{2}+\frac{(p(\xi))^{2}}{mn(\xi)},\quad U=U_{0}+\frac{1}{\nu},$ (9)
where $n$ and $p$ are particle and momentum densities. After these rearrangements, transport equations read
$\left(\gamma(n)+\partial_{t}\right){\boldsymbol{p}}(\xi)+\partial_{{\boldsymbol{x}}}\Pi(\xi)=en{\boldsymbol{E}}$ (10)
$\partial_{t}n(\xi)+\frac{1}{m}\partial_{{\boldsymbol{x}}}{\boldsymbol{p}}(\xi)=0,$ (11)
where in the continuity equation we replaced $nv$ with $p/m$. A steady uniform current-carrying state obeys
$\gamma{\boldsymbol{p}}=en{\boldsymbol{E}},$ (12)
yielding a standard result for the flow velocity, $\gamma u=\frac{e}{m}E$.
Next, we show that Eqs. (10) and (11) predict instability of a current-carrying state. In performing numerics, it is beneficial to regularize the dynamics at large wavenumbers by introducing diffusion-like terms in the $p$ and $n$ equations. This can be done by replacing
$\partial_{t}p\to(\partial_{t}-D_{p}\partial_{x}^{2})p,\quad\partial_{t}n\to(\partial_{t}-D_{n}\partial_{x}^{2})n$ (13)
in Eqs.(10) and (11). This change does not affect the behavior at long wavelengths, provided $D_{n}$ and $D_{p}$ are small, however it makes the dynamics well behaved at short wavelengths, allowing to perform numerical simulations and explore the behavior at long times where the nonlinear behavior triggered by Turing instability emerges—see Figs. 1 and 2, and discussion in [41].
Next, we examine the conditions under which the transport equations develop an instability. Perturbing the steady state as $n=\bar{n}+\delta n$, $p=\bar{p}+\delta p$, and linearizing Eqs.(10) and (11) in $n$ and $p$, gives
$\left(\gamma_{n}\bar{p}-eE\right)\delta n+\Pi_{n}\partial_{x}\delta n+\gamma\delta p+(\partial_{t}-D_{p}\partial_{x}^{2})\delta p$
$+\Pi_{p}\partial_{x}\delta p=0,\quad(\partial_{t}-D_{n}\partial_{x}^{2})\delta n+\frac{1}{m}\partial_{x}\delta p=0,$ (14)
where the quantities with subscripts $n$ and $p$ denote partial derivatives: $\gamma_{n}=\partial\gamma/\partial n$, $\Pi_{n}=\partial\Pi/\partial n$ and $\Pi_{p}=\partial\Pi/\partial p$. For harmonic perturbations $\delta n(\xi)=\delta n_{\omega,k}e^{ikx-i\omega t}$, $\delta p(\xi)=\delta p_{\omega,k}e^{ikx-i\omega t}$ this problem yields a pair of coupled linear equations
$(\Lambda+ik\Pi_{n})\delta n_{\omega,k}$ $+(\gamma-i\omega+D_{p}k^{2}+ik\Pi_{p})\delta p_{\omega,k}=0$
$(-i\omega+D_{n}k^{2})$ $\delta n_{\omega,k}+i\frac{k}{m}\delta p_{\omega,k}=0$ (15)
where we used a shorthand notation $\Lambda=\left(\frac{\partial\gamma}{\partial n}-\frac{\gamma}{n}\right)p$. Instability occurs when the linearized problem admits a complex-frequency solution growing with time,
$\delta n_{k}(t),\,\delta p_{k}(t)\sim e^{-i\omega t},\quad{\rm Im}\,\omega_{+}(k)>0.$ (16)
The characteristic equation for this problem, in matrix form, reads
$\left(\begin{array}[]{cc}\Lambda+ik\Pi_{n}&\gamma-i\omega+D_{p}k^{2}+ik\Pi_{p}\\ -i\omega+D_{n}k^{2}&i\frac{k}{m}\end{array}\right)\left(\begin{array}[]{c}\delta n_{\omega,k}\\ \delta p_{\omega,k}\end{array}\right)=0.$ (17)
This yields two normal modes with frequencies
$\omega_{\pm}(k)=\frac{-i\tilde{\gamma}+k\Pi_{p}\pm i\sqrt{\Delta}}{2}-iD_{n}k^{2},$ (18)
where $\tilde{\gamma}=\gamma+(D_{p}-D_{n})k^{2}$ and
$\Delta=\left(\tilde{\gamma}+ik\Pi_{p}\right)^{2}+\frac{4ik}{m}\Lambda-\frac{4k^{2}}{m}\Pi_{n}.$ (19)
This general result is applicable to systems with both low and high dissipation, and both in the absence and in the presence of the flow. At $k=0$ the frequency $\omega_{+}$ vanishes, whereas the frequency $\omega_{-}$ takes a negative-imaginary value, corresponding to the conserved density and relaxing velocity modes, respectively.
Before analyzing amplification and instability in the strongly dissipative regime $\gamma\gg\omega$—the main focus of this work—it is useful to verify that, in the weak-dissipation limit, our results reproduce the known plasma mode dispersions, both with and without steady background flow. Setting $\gamma$, $\Lambda$, $D_{n}$ and $D_{p}$ to zero gives
$\omega_{\pm}(k)=\frac{1}{2}k\Pi_{p}\pm\sqrt{\frac{1}{4}k^{2}\Pi_{p}^{2}+\frac{1}{m}k^{2}\Pi_{n}},$ (20)
where we moved $i$ under the square root. Focusing on a parabolic band and plugging $\Pi_{p}=2p/nm$ and $\Pi_{n}=Un-p^{2}/n^{2}m$ [with $U$ the sum of the ee interaction and Fermi sea pressure contributions, as given in Eq.(9)] the $p$ dependence under the square root cancels, giving
$\omega_{\pm}(k)=ku\pm ku_{0},\quad u_{0}=\sqrt{Un/m}.$ (21)
Here $u_{0}$ is plasmon velocity at density $n$ and $u$ is the flow velocity, $p=mnu$. This gives the Doppler-shifted undamped plasma dispersion $(\omega-uk)^{2}=k^{2}u_{0}^{2}$, as expected from Galilean symmetry for a parabolic band.
Next, we return to the analysis of dissipative transport and estimate maximally unstable wavelength and instability threshold. It is instructive to focus on the strongly dissipative regime, where the momentum relaxation rate $\gamma$ greatly exceeds scales such as $\omega p$ and $k\Pi$. For the parabolic band model the general expression for $\omega_{\pm}$ given in Eqs. (18) and (19) simplifies considerably. Substituting the derivatives $\Pi_{n}$ and $\Pi_{p}$ in Eq. (19) yields
$\Delta=\gamma^{2}+ik\frac{4p}{m}\frac{d\gamma}{dn}-k^{2}\frac{4Un}{m}$ (22)
An analytic estimate of the most unstable wavelength $k_{*}$ can be obtained by examining the instability growth rate ${\rm Im}\,\omega_{+}(k)$ vs. $k$ in the limits of small and large $k$, and matching the two asymptotic behaviors in the intermediate (“borderline”) region. For clarity, we focus on the case of small diffusion coefficients $D_{n}$ and $D_{p}$, setting $D_{n}=D_{p}=0$. In the large-$k$ limit ($k\gg k_{*}$), the growth rate approaches a $k$-independent value:
${\rm Im}\,\omega_{+}(k)_{k\gg k_{*}}=-\frac{1}{2}\gamma+\frac{s}{2u_{0}}\frac{pd\gamma}{mdn},\quad s={\rm sgn}\,\left(p\frac{d\gamma}{dn}\right).$ (23)
This predicts the instability threshold at
$p\frac{d\gamma}{dn}=s\gamma mu_{0}$ (24)
with an unstable phase for $p\frac{d\gamma}{dn}s>\gamma mu_{0}$ and a stable phase for $p\frac{d\gamma}{dn}<\gamma mu_{0}$. Namely, the instability occurs when the flow velocity $u=p/mn$ exceeds the critical value $u_{c}$ given in Eq.(2). If $R$ is order-one, the flow velocity at which instability occurs is comparable to the plasmonic velocity $u_{0}=\sqrt{Un/m}$.
Next, to complete the estimate of the most unstable wavelength we consider small $k$, where Taylor expansion of ${\rm Im}\,\omega_{+}(k)$ to leading order in $k$ gives
${\rm Im}\,\omega_{+}(k)_{k\ll k_{*}}=k^{2}\left[\frac{p^{2}\left(\frac{d\gamma}{dn}\right)^{2}}{m^{2}\gamma^{3}}-\frac{u_{0}^{2}}{\gamma}\right]+O(k^{4})$ (25)
For $k\approx k_{*}$, these values should match those in Eq.(23), obtained at $k\gg k_{*}$. Writing this condition as
$k_{*}^{2}\left[\frac{p^{2}\left(\frac{d\gamma}{dn}\right)^{2}}{m^{2}\gamma^{3}}-\frac{u_{0}^{2}}{\gamma}\right]=\frac{p\frac{d\gamma}{dn}}{2mu_{0}}-\frac{1}{2}\gamma$ (26)
and noting that both the left- and the right-hand side vanish at the transition but remain nonzero away from transition, we obtain [51]:
$k_{*}^{2}\approx\frac{1}{4}\frac{\gamma^{2}}{u_{0}^{2}}.$ (27)
Notably, the maximally unstable wavelength, $\lambda_{*}=2\pi/k_{*}$, remains finite at the transition. This behavior is consistent with our simulation results and mirrors the finite modulation wavelength observed at the onset of the hydrodynamic Kapitsa instability[34, 35]. Density-dependent $\gamma$ and $u_{0}$ lead to tunable drift velocity $u_{d}$ and wavelength spanning a range of values, from submicron to hundreds of microns. Accordingly, the narrow-band radiation frequency, given by the washboard formula $f=u_{d}/\lambda$, for realistic system parameters can vary from sub-GHz to hundreds of GHz. For estimates and plots of critical velocity and modulation wavelength see [41].
In summary, we identify an electron–hydrodynamic instability in Dirac materials that is closely analogous to the Turing-Kapitsa instability generating trains of roll waves in viscous films. In electron systems, strong density dependence of current dissipation near charge neutrality makes the flow susceptible to modulation, triggering a transition at a critical drift velocity $u_{c}$, beyond which the system develops self-sustained running waves. This instability produces both a sharp, second-order–like change in the time-averaged current and narrow-band emission at the washboard frequency $f=u/\lambda$. Estimated frequencies span a wide, current-tunable range extending to hundreds of GHz, highlighting Dirac materials as a promising platform for high-frequency electron-fluid phenomena.
This work greatly benefited from discussions with Ray Ashoori, Serguei Brazovskii, Andre Geim, Frank Koppens, Brian Skinner and Eli Zeldov.