### v2

# thanks: Equal contributionsthanks: Equal contributions

Abstract
We predict hydrodynamic Turing instability of current-carrying Dirac electron fluids that drives spontaneous self-oscillatory transport. The instability arises near charge neutrality, where carrier kinetics make current dissipation strongly density dependent. Above a critical drift velocity, a uniform electronic flow becomes unstable and undergoes a dynamical transition to a state with coupled spatial modulation and temporal oscillations—an electronic analogue of Kapitsa roll waves in viscous films. The transition exhibits two clear signatures: a nonanalytic, second-order–like onset in the time-averaged current and narrow-band electromagnetic emission at a tunable washboard frequency $f=u/\lambda$. Although reminiscent of sliding charge-density waves, the mechanism is intrinsic and disorder independent. Owing to the small effective mass of Dirac carriers, hydrodynamic time scales translate into emission frequencies in the tens to hundreds of gigahertz range, establishing Dirac materials as a platform for high-frequency self-oscillatory electron hydrodynamics.
Pattern formation is a hallmark of driven many-body systems, revealing how collective behavior can emerge even without external structuring. A paradigmatic example is the Turing instability [1, 2, 3]. Despite its universality, this mechanism has been explored almost exclusively in classical settings. Here we predict a quantum analogue: a modulational (Turing-type) instability in quantum materials driven out of equilibrium by a DC current. The key idea is that an applied current selectively softens the system at particular wave numbers, triggering an instability toward spontaneous spatial modulation. We find that current-driven Turing effects naturally arise in quantum dissipative systems, such as Dirac and Weyl metals in which electron-hole scattering enhances resistivity near charge neutrality while remaining weak away from it [4, 5, 6, 7, 8, 9, 10, 11].
Beyond fundamental interest, determining when a current-carrying state becomes unstable to self-oscillation may also clarify a longstanding practical question: Can a DC current generate coherent high-frequency radiation? Proposed mechanisms have ranged from Bloch oscillations [12, 13, 14, 15] to plasmonic instabilities [16, 17], and while experimental signatures exist [18, 19, 20, 21, 22], achieving robust and tunable self-oscillation has remained difficult. Our results identify a new route to current-driven pattern formation and self-oscillation— arising from the interplay of dissipation, band geometry, and hydrodynamic effects in Dirac materials — and highlight fresh opportunities for exploring quantum systems with complex band geometry or strong correlations.
Figure 1: (a) Current dissipation in graphene bilayer (BLG) band vs. electron density. Dissipation is high in the Dirac plasma near charge neutrality point (CNP) but drops rapidly upon doping away from CNP. Density-dependent dissipation drives the Turing–Kapitsa instability. Insets illustrate carrier population in BLG band for densities at CNP and away from it, respectively. (b) Instability growth rate ${\rm Im}\,\omega_{\pm}(k)$ for the modes $\omega_{+}(k)$ and $\omega_{-}(k)$, Eq. (19), representing density and velocity perturbations. The density mode satisfies $\omega_{+}(0)=0$ by particle-number conservation, becoming unstable above a critical flow velocity (Eq.(2)); the growth rate ${\rm Im}\,\omega_{+}(k)>0$ peaks at the wave numbers $\pm k_{*}$ marked by arrows. The velocity mode remains damped, ${\rm Im}\,\omega_{-}(k)<0$. Small initial perturbations grow into a periodic modulation with wavelength $\lambda\sim 2\pi/k_{*}$ propagating downstream, as shown in Fig. 2.
The instability mechanism proposed here is reminiscent of the Kapitsa instability in viscous films flowing down an incline [23, 24, 25, 26, 27, 28]. The emergence of periodic modulations in steady laminar flows of open-surface films is a direct consequence of the balance between gravity and viscous drag: $\rho g\sin\theta=\eta\nabla^{2}{\boldsymbol{v}}({\boldsymbol{r}})$, where $\theta$ is the incline angle, $g$ gravity, and $\eta$ viscosity. A steady state defines a scaling between flow velocity and film thickness, $u\propto w^{2}g/\eta$. Since $u$ grows rapidly with $w$, regions with slightly larger film thickness $w(x)$ will flow faster, thereby amplifying perturbations and giving rise to a self-sustained running wave,
$w(x)=\bar{w}+\delta w\cos k(x-ut).$ (1)
In fluid films, the wavelength $\lambda=2\pi/k$ and modulation amplitude are set by the competition between viscous and capillary forces [24]. The emerging wavenumber $k$ coincides with $k_{*}$—the wavevector at which the instability growth rate is maximal—mirroring the essence of Turing’s pattern-formation mechanism [25].
As we will see, an electronic analogue of this instability can cause a steady (DC) electron flow to spontaneously develop self-sustained charge and current oscillations. The correspondence between an atomically thin 2D conductor with carrier density $n(x,y)$ and a classical viscous film with spatially varying thickness w(x,y)w(x,y)—measured between the substrate and its free surface—is established by identifying $n(x,y)$ with $w(x,y)$. Beyond this “hyperspace mapping” $w(x,y)\to(x,y,n)$, realizing the Turing–Kapitsa instability requires a dissipation mechanism whose rate, being strong, varies rapidly with nn—the electronic analogue of a drag coefficient in viscous films that decreases sharply with increasing film thickness.
Figure 2: Wavelength selection at instability. Shown is the long-time modulation (orange) evolved from different initial perturbations (blue): $\delta n(x)=\sum_{m}a_{m}\cos k_{m}x+b_{m}\sin k_{m}x$, $\delta p(x)=\sum_{m}a^{\prime}_{m}\cos k_{m}x+b^{\prime}_{m}\sin k_{m}x$, with $k_{m}=\frac{2\pi}{L}\times 3,5,8,13,21,55$. Irrespective of initial conditions, the system converges to the same spatial pattern, with wavelength set by the wavenumber $k$ of maximal growth (marked by arrows in Fig. 1), consistent with Turing’s maximum-growth rule.
Several effects reported in the literature can cause dissipation to decrease as carrier concentration increases. Here we focus on quasiparticle scattering at charge neutrality (CNP) [4, 5, 6, 7, 8, 9, 10, 11]. This dissipation channel is strong at CNP, where the Dirac plasma is e–h balanced, but it weakens rapidly once the carrier density is tuned away from neutrality. The resulting density-dependent dissipation plays a role similar to the width-dependent viscous drag in thin-film hydrodynamics: when the drift velocity exceeds a threshold, the flow becomes unstable and develops downstream-propagating waves.
Our analysis predicts a critical drift velocity:
$u_{c}=\frac{u_{0}}{R},\qquad R=\frac{n}{\gamma}\left|\frac{d\gamma}{dn}\right|,\qquad u_{0}=\sqrt{\frac{Un}{m}},$ (2)
where the dimensionless quantity $R$ measures how sensitively the dissipation rate $\gamma$ responds to changes in carrier density, while $u_{0}$ denotes the 2D plasmon velocity at density $n$. Because the dissipative regime of interest lies close to the charge neutrality point (CNP), the relevant carrier densities are much smaller—i.e., much closer to neutrality—than those typically used in studies of ballistic 2D plasmon propagation [29, 30, 31]. As shown in Fig. 1(a), the strong temperature dependence of $\gamma$ produces a pronounced density dependence of $R$ near the CNP. As a result, the critical velocity $u_{c}$ varies strongly with both temperature and carrier density [32].
For $u>u_{c}$, the system develops a self-sustained modulation that follows Turing’s wavelength-selection rule. Close to threshold, the modulation wavelength is $\lambda=2\pi/k_{*}$, where $k_{*}$ is the most unstable mode, at which the instability growth rate is maximal (see Figs. 1(b) and 2). The transition is continuous and second-order (Fig. 3), and the modulation wavelength remains finite at the threshold, as in Brazovskii and Swift-Hohenberg theories of convective instabilities[33, 34], without showing any divergence. For realistic parameters, $\lambda$ spans a wide range, from submicron to hundreds of microns [32].
Observable signatures of this collective dynamics resemble those studied for other running wave phenomena, such as sliding charge-density and spin-density waves (CDW and SDW) [35, 36, 37]. One is a narrow-band emission at the “washboard” frequency $f=u/\lambda$, where $u$ is the flow velocity tunable by the applied DC current. Another is a characteristic nonlinearity in the current–field dependence at the instability threshold, shown in Fig.3. The increase in the slope of the current vs. $u$ dependence reflects the contribution of running waves to nonlinear conductivity, resembling the Lee-Fukuyama-Rice conductivity of sliding CDW[35, 36, 37].
We stress that both the AC emission by running waves at a frequency $f=u/\lambda$ and the DC current nonlinearity are intrinsic effects, present already in a homogeneous system. This is in contrast to CDW sliding, where narrow-band emission and current nonlinearity are governed by the effects due to pinning-depinning by disorder.
Figure 3: Signatures and observables for current-driven electronic Turing–Kapitsa instability. Plotted are the running waves modulation amplitude and time-averaged current near the instability onset, obtained numerically for a graphene bilayer model (see Eqs. (11), (12) and accompanying text). The red line shows the theoretical square-root behavior $\Delta n\propto(u_{d}-u_{d}^{*})^{1/2}$ expected on the general symmetry grounds [45]. The sharp slope change in the time-averaged (DC) current–field dependence marks the instability threshold and the added current from the running waves. Insets (a) and (b) illustrate a typical modulation and the current–field relation.
To explore the emergence of running waves under Turing-Kapitsa instability driven by a DC current, we consider a simple but realistic transport model. A natural framework to analyze collective dynamics of electrons is provided by hydrodynamic equations for particle number and momentum density $n(x,t)$ and ${\boldsymbol{p}}(x,t)$. Microscopically, these quantities are
$n(x,t)=\sum_{{\boldsymbol{p}}^{\prime}}f({\boldsymbol{p}}^{\prime},x,t),\quad{\boldsymbol{p}}(x,t)=\sum_{{\boldsymbol{p}}^{\prime}}{\boldsymbol{p}}^{\prime}f({\boldsymbol{p}}^{\prime},x,t),$ (3)
where $f({\boldsymbol{p}},x,t)$ is particle phase-space distribution summed over spins and valleys that evolves in time under the kinetic equation [38, 39]
$(\partial_{t}+{\boldsymbol{v}}\cdot\partial_{{\boldsymbol{x}}}+{\boldsymbol{F}}\cdot\partial_{{\boldsymbol{p}}})f({\boldsymbol{p}},\xi)=I[f].$ (4)
Here, ${\boldsymbol{v}}=\partial_{{\boldsymbol{p}}}\epsilon(p)$ is the velocity, ${\boldsymbol{F}}=e{\boldsymbol{E}}$ the force, and $I[f]$ the collision integral describing electron-electron and electron-hole scattering, and interactions with disorder and phonons; $\xi=(x,t)$ denotes space and time variables.
We are interested in transport in a Dirac band of a graphene bilayer, $\varepsilon_{\pm}({\boldsymbol{k}})=\pm\tfrac{1}{2}\!\left(\sqrt{\gamma_{1}^{2}+4(\hbar vk)^{2}}-\gamma_{1}\right)$, where $v\approx 10^{6}\,$m/s and $\gamma_{1}=0.39\,$eV. For simplicity, we consider a one-species model representing one component of the bilayer bands—either electron or hole. The results below are largely insensitive to the exact form of $\epsilon(p)$ or to microscopic details of how momentum dissipation varies with carrier density. We model this by a dissipation rate that peaks in the region of width $w\sim k_{B}T$ and decreases with detuning from the CNP[10, 11]:
$\gamma(n)=\gamma_{0}e^{-|n|/w},\quad w=\frac{2m}{\hbar^{2}}k_{B}T,\quad m\approx 0.04m_{\rm e},$ (5)
where $w$ describes the CNP width due to intrinsic e-h scattering in pristine graphene in a parabolic band approximation. In this model, the electron temperature is replaced by an effective time-averaged value, neglecting its dynamics. Although heating is important, it occurs on time scales much longer than those of oscillatory electron motion. Moreover, recent experiments[40, 41, 42] show that high currents can flow through moiré graphene at drift velocities approaching the band velocity without significant overheating. Efficient cooling via the hBN substrate enables such high-current operation without damaging the graphene or broadening the Dirac band at CNP.
When applied to the particle number and momentum densities, $n$ and $p$, Eq.(4) gives conservation laws:
$\left(\gamma(n)+\partial_{t}\right)$ ${\boldsymbol{p}}(\xi)+\partial_{{\boldsymbol{x}}}\Pi_{0}(\xi)=en\left({\boldsymbol{E}}-\partial_{{\boldsymbol{x}}}\Phi(\xi)\right)$ (6)
$\partial_{t}$ $n(\xi)+\partial_{{\boldsymbol{x}}}[n(\xi){\boldsymbol{v}}(\xi)]=0.$ (7)
Here the first equation is the continuum version of the second Newton’s law $dp/dt=F$, the second equation is the continuity equation for $n$. In a parabolic band, momentum density is related to particle current density ${\boldsymbol{j}}=\sum_{p^{\prime}}{\boldsymbol{v}}(p^{\prime})f({\boldsymbol{p}}^{\prime})\equiv n{\boldsymbol{v}}$ as $m{\boldsymbol{j}}={\boldsymbol{p}}$.
The quantity $\Pi_{0}$ is the Fermi pressure of a flowing Fermi sea described by Doppler-shifted Fermi distribution $f({\boldsymbol{p}})=\frac{1}{e^{\beta(\epsilon(p)-{\boldsymbol{u}}\cdot{\boldsymbol{p}}-\mu)}+1}$. For electron gas of density $n$ flowing with the velocity $u$, for parabolic band, $\Pi_{0}$ is a sum of a static contribution and a kinetic energy term
$\Pi_{0}=\sum_{p^{\prime}}p_{x}^{\prime}v_{x}(p^{\prime})f({\boldsymbol{p}}^{\prime})=\frac{n^{2}}{2\nu}+\frac{p^{2}}{nm},\quad p=mnu,$ (8)
where $\nu=n/\epsilon_{F}=gm/2\pi\hbar^{2}$ is the density of states ($g$ is spin and valley degeneracy). The simple form of the result in Eq.(8), which is a sum of the pressure of a Fermi sea at rest $\Pi_{0}(n)=\sum_{p^{\prime}}\epsilon(p^{\prime})f_{u=0}(p^{\prime})=n^{2}/2\nu$ and the Bernoulli pressure $p^{2}/nm$ due to the flow, is a property of a parabolic band arising from its Galilean symmetry.
The transport equations (6) and (7) capture the essential physics underlying the current-driven Turing instability. A more realistic description, however, must explicitly include the interaction and viscous effects incorporated in the hydrodynamic stress density $\Pi$.
Below, for conciseness, we use a constant interaction model deferring the discussion of long-range interaction till later. In this case, the last term in (6) takes the form
$\Phi(\xi)=U_{0}[n(\xi)-\bar{n}],$ (9)
where constant $\bar{n}$ represents an offset charge on a gate.
It is convenient to combine the terms $\partial_{{\boldsymbol{x}}}\Pi_{0}$ and $\partial_{{\boldsymbol{x}}}\Phi$ in Eq. (6) into a net Fermi pressure term $\partial_{{\boldsymbol{x}}}\Pi$ that accounts for both electron–electron interactions and the flow, [43]
$\Pi=\frac{1}{2}U\left(n(\xi)\right)^{2}+\frac{(p(\xi))^{2}}{mn(\xi)},\quad U=U_{0}+\frac{1}{\nu},$ (10)
where $n$ and $p$ are particle and momentum densities. After these rearrangements, transport equations read
$\left(\gamma(n)+\partial_{t}\right){\boldsymbol{p}}(\xi)+\partial_{{\boldsymbol{x}}}\Pi(\xi)=en{\boldsymbol{E}}$ (11)
$\partial_{t}n(\xi)+\frac{1}{m}\partial_{{\boldsymbol{x}}}{\boldsymbol{p}}(\xi)=0,$ (12)
where in the continuity equation we replaced $nv$ with $p/m$. A steady uniform current-carrying state obeys
$\gamma{\boldsymbol{p}}=en{\boldsymbol{E}},$ (13)
yielding a standard result for the flow velocity, $\gamma u=\frac{e}{m}E$.
Next, we show that Eqs. (11) and (12) predict instability of a current-carrying state. In performing numerics, it is beneficial to regularize the dynamics at large wavenumbers by introducing diffusion-like terms in the $p$ and $n$ equations. This can be done by replacing
$\partial_{t}p\to(\partial_{t}-D_{p}\partial_{x}^{2})p,\quad\partial_{t}n\to(\partial_{t}-D_{n}\partial_{x}^{2})n$ (14)
in Eqs.(11) and (12). This change does not affect the behavior at long wavelengths, provided $D_{n}$ and $D_{p}$ are small, however it makes the dynamics well behaved at short wavelengths, allowing to perform numerical simulations and explore the behavior at long times where the nonlinear behavior triggered by Turing instability emerges—see Figs. 1 and 2, and discussion in [32].
Next, we examine the conditions under which the transport equations develop an instability. Perturbing the steady state as $n=\bar{n}+\delta n$, $p=\bar{p}+\delta p$, and linearizing Eqs.(11) and (12) in $n$ and $p$, gives
$\left(\gamma_{n}\bar{p}-eE\right)\delta n+\Pi_{n}\partial_{x}\delta n+\gamma\delta p+(\partial_{t}-D_{p}\partial_{x}^{2})\delta p$
$+\Pi_{p}\partial_{x}\delta p=0,\quad(\partial_{t}-D_{n}\partial_{x}^{2})\delta n+\frac{1}{m}\partial_{x}\delta p=0,$ (15)
where the quantities with subscripts $n$ and $p$ denote partial derivatives: $\gamma_{n}=\partial\gamma/\partial n$, $\Pi_{n}=\partial\Pi/\partial n$ and $\Pi_{p}=\partial\Pi/\partial p$. For harmonic perturbations $\delta n(\xi)=\delta n_{\omega,k}e^{ikx-i\omega t}$, $\delta p(\xi)=\delta p_{\omega,k}e^{ikx-i\omega t}$ this problem yields a pair of coupled linear equations
$(\Lambda+ik\Pi_{n})\delta n_{\omega,k}$ $+(\gamma-i\omega+D_{p}k^{2}+ik\Pi_{p})\delta p_{\omega,k}=0$
$(-i\omega+D_{n}k^{2})$ $\delta n_{\omega,k}+i\frac{k}{m}\delta p_{\omega,k}=0$ (16)
where we used a shorthand notation $\Lambda=\left(\frac{\partial\gamma}{\partial n}-\frac{\gamma}{n}\right)p$. Instability occurs when the linearized problem admits a complex-frequency solution growing with time,
$\delta n_{k}(t),\,\delta p_{k}(t)\sim e^{-i\omega t},\quad{\rm Im}\,\omega_{+}(k)>0.$ (17)
The characteristic equation for this problem, in matrix form, reads
$\left(\begin{array}[]{cc}\Lambda+ik\Pi_{n}&\gamma-i\omega+D_{p}k^{2}+ik\Pi_{p}\\ -i\omega+D_{n}k^{2}&i\frac{k}{m}\end{array}\right)\left(\begin{array}[]{c}\delta n_{\omega,k}\\ \delta p_{\omega,k}\end{array}\right)=0.$ (18)
This yields two normal modes with frequencies
$\omega_{\pm}(k)=\frac{-i\tilde{\gamma}+k\Pi_{p}\pm i\sqrt{\Delta}}{2}-iD_{n}k^{2},$ (19)
where $\tilde{\gamma}=\gamma+(D_{p}-D_{n})k^{2}$ and
$\Delta=\left(\tilde{\gamma}+ik\Pi_{p}\right)^{2}+\frac{4ik}{m}\Lambda-\frac{4k^{2}}{m}\Pi_{n}.$ (20)
This general result is applicable to systems with both low and high dissipation, and both in the absence and in the presence of the flow. At $k=0$ the frequency $\omega_{+}$ vanishes, whereas the frequency $\omega_{-}$ takes a negative-imaginary value, corresponding to the conserved density and relaxing velocity modes, respectively.
Figure 4: Modulation wavelength $\lambda=2\pi/k_{*}$, where $\pm k_{*}$ (blue arrows) maximizes the instability growth rate $\mathrm{Im}\,\omega_{+}(k)$, is governed by the behavior of $\mathrm{Im}\,\omega_{+}(k)$ at small and large $k$. For small $k$, the growth rate (red parabola) behaves as $\alpha k^{2}$, with $\alpha>0$ for $u>u_{c}$ and $\alpha<0$ for $u<u_{c}$. For large $k$, it saturates to a $k$-independent value (horizontal red line), again positive for $u>u_{c}$ and negative for $u<u_{c}$. Matching these small- and large-$k$ asymptotics yields Eq. (24), and hence Eq. (26). The instability growth rate is shown for $u=0.55>u_{c}$ and $u=0.2<u_{c}$ (blue and pink curves); solid and dashed lines correspond to $D_{n,p}=0.1$ and $D_{n,p}=0$.
Prior to analyzing the dissipative regime $\gamma\gg\omega$—the main focus of this work—we verify that, in the weak-dissipation limit, our results reproduce the known plasma mode dispersions, both with and without steady background flow. Setting $\gamma$, $\Lambda$, $D_{n}$ and $D_{p}$ to zero gives
$\omega_{\pm}(k)=\frac{1}{2}k\Pi_{p}\pm\sqrt{\frac{1}{4}k^{2}\Pi_{p}^{2}+\frac{1}{m}k^{2}\Pi_{n}},$ (21)
where we moved $i$ under the square root. Focusing on a parabolic band and plugging $\Pi_{p}=2p/nm$ and $\Pi_{n}=Un-p^{2}/n^{2}m$ [with $U$ the sum of the ee interaction and Fermi sea pressure contributions, as given in Eq.(10)] the $p$ dependence under the square root cancels, giving
$\omega_{\pm}(k)=ku\pm ku_{0},\quad u_{0}=\sqrt{Un/m}.$ (22)
Here $u_{0}$ is plasmon velocity at density $n$ and $u$ is the flow velocity, $p=mnu$. This gives the Doppler-shifted undamped plasma dispersion $(\omega-uk)^{2}=k^{2}u_{0}^{2}$, as expected from Galilean symmetry for a parabolic band.
Next, we estimate maximally unstable wavelength and instability threshold for the dissipative regime, where the relaxation rate $\gamma$ greatly exceeds scales such as $\omega p$ and $k\Pi$. For a parabolic band model the general expression for $\omega_{\pm}$ given in Eqs. (19) and (20) simplifies considerably. Substituting the derivatives $\Pi_{n}$ and $\Pi_{p}$ in Eq. (20) yields
$\Delta=\gamma^{2}+ik\frac{4p}{m}\frac{d\gamma}{dn}-4k^{2}u_{0}^{2}$ (23)
An analytic estimate of the most unstable wavenumber $k_{\ast}$ can be obtained as illustrated in Fig. 4. We compare the instability growth rate ${\rm Im}\,\omega_{+}(k)$ in the small- and large-$k$ limits and match these asymptotic behaviors in an intermediate (“borderline”) region. For clarity, we focus on the case of small $D_{n}$ and $D_{p}$. In the large- and small-$k$ limits, the growth rate behaves as
${\rm Im}\,\omega_{+}(k)=\begin{cases}-\frac{1}{2}\gamma+\frac{\left|p\frac{d\gamma}{dn}\right|}{2mu_{0}},&k\gg k_{*},\\ k^{2}\frac{\left[p^{2}\left(\frac{d\gamma}{dn}\right)^{2}-m^{2}\gamma^{2}u_{0}^{2}\right]}{m^{2}\gamma^{3}}&k\ll k_{*}.\end{cases}$ (24)
The sign change predicts instability threshold $u=u_{c}$ with the unstable and stable phases for $\left|p\frac{d\gamma}{dn}\right|-\gamma mu_{0}$ positive and negative, respectively. The sign change occurs simultaneously for $k\gg k_{*}$ and $k\ll k_{*}$. This gives critical velocity $u_{c}$ given in Eq.(2). If $R\sim 1$, the critical velocity is comparable to the plasmonic velocity $u_{0}$ in a dissipation-free system at density close to CNP.
An estimate of the most unstable mode $k_{\star}$ follows from matching the small- and large-$k$ behavior Eq.(24):
$\frac{k_{*}^{2}}{m^{2}\gamma^{3}}\left[p^{2}\left(\frac{d\gamma}{dn}\right)^{2}-m^{2}\gamma^{2}u_{0}^{2}\right]=\frac{\left|p\frac{d\gamma}{dn}\right|}{2mu_{0}}-\frac{1}{2}\gamma.$ (25)
Solving for $k_{*}$ and noting that both sides vanish at the transition but remain nonzero away from transition, in the limit $\left|p\frac{d\gamma}{dn}\right|\to mu_{0}\gamma$ we obtain [44]:
$k_{*}^{2}\approx\frac{1}{4}\frac{\gamma^{2}}{u_{0}^{2}}.$ (26)
Notably, the maximally unstable wavelength $\lambda_{\ast}=2\pi/k_{\ast}$ remains finite at the transition, in agreement with our simulations and in analogy with the finite modulation wavelength at the onset of the hydrodynamic Kapitsa instability.[25, 26] Density-dependent $\gamma$ and $u_{0}$ yield a tunable critical velocity $u_{c}$ and wavelength ranging from submicron to hundreds of microns. The resulting narrow-band radiation frequency, $f=u_{d}/\lambda$, can thus span from sub-GHz to hundreds of GHz for realistic parameters; see [32] for estimates and plots of the critical velocity and modulation wavelength.
In summary, we identify an electron–hydrodynamic instability in Dirac materials closely analogous to the Turing–Kapitsa instability that generates roll-wave trains in viscous films. Density dependence of current dissipation near charge neutrality makes the flow unstable above a critical drift velocity $u_{c}$, resulting in self-sustained running waves. This leads to a sharp, second-order–like change in the time-averaged current and narrow-band emission at the washboard frequency $f=u/\lambda$, with a current-tunable range extending to hundreds of GHz, highlighting Dirac materials as a promising platform for high-frequency electron-fluid phenomena.
This work greatly benefited from discussions with Ray Ashoori, Serguei Brazovskii, Andre Geim, Frank Koppens, Leo Radzihovsky, Brian Skinner, and Eli Zeldov.